Check whether the function: f: R^2→R, defined by

f(x,y)= x+ysinx has extremum at any point in the domain of f.



1
Expert's answer
2021-04-26T08:25:28-0400

"\bigstar"


The function is defined in domain




"\boxed{f(x,y)= x+ysinx}" .......(1)


"\bigstar" solution

=

"\frac{\partial x+ysin(x)}{\partial x}=0\\ \frac{\partial (x+ysin(x))}{ \partial y}=0" ..................(2)



In equation (1) we do partial differentiation

w.r.t both x and y



w.r.t x we get


1 + ycos(x)= 0...............(3)

And


w.r.t y


0 + sin(x) = 0..........(4)


And using different equations (1),(2),(3),(4)



we got



"\boxed{x = 2n\pi}" and "\boxed{y= -1}" and "n\in z"


"\boxed{x = 2n\pi + \pi}" and "\boxed{y= 1}" and n"\in" z


On partial differentiating w.r.t x and y

we got above values.



for better understand

we have different graphs


3D graph

For the function

"\boxed{f(x,y)= x+ysinx}"














CONTOUR PLOT for the function

"\boxed{f(x,y)= x+ysinx}"